# Constraint Breeds Generalization: Temporal Dynamics as an Inductive Bias

### v1

@@ A.1 Experiment 1: Cross-Encoding Generalization Details
Dataset and Setup. We use the sklearn.datasets.load_digits dataset (64 features, 10 classes)[optical_recognition_of_handwritten_digits_80, scikit-learn], split into 70% training, 15% validation, and 15% testing. Input features $\mathbf{x}$ are first normalized using StandardScaler. Networks are trained on data encoded with one $\delta_{\text{train}}$ value, then tested on 12 different encodings: $\delta_{\text{test}}\in\{-1.5,-1.0,-0.3,0.0,0.3,1.0,1.5,2.0,2.5,5.0,7.0,10.0\}$. The detailed quantitative generalization matrices (Mean $\pm$ Std %) for all architectures are provided in Tables 5 - 9.
Training. LIF neurons use learnable thresholds and membrane decay $\beta=0.95$. Training uses surrogate gradients (surrogate.fast_sigmoid with slope=25)[eshraghian2023training] with the Adam optimizer (lr=1e-4)[kingma2014adam] for 200 epochs, with early stopping (patience=10) based on validation accuracy.
@@ A.2 Experiment 2: Unsupervised Receptive Field Learning Details
Dataset and Setup. The model for this experiment is a spiking neural network (SNN) autoencoder. Its architecture consists of an encoder linear layer ($W_{\text{enc}}\in\mathbb{R}^{128\times d}$), a recurrent Leaky Integrate-and-Fire (LIF) hidden layer with 128 learnable thresholds and membrane decay $\beta=0.95$ neurons[eshraghian2023training], and a decoder linear layer ($W_{\text{dec}}\in\mathbb{R}^{d\times 128}$). The sparse autoencoder minimizes a composite loss balancing reconstruction fidelity and code sparsity, inspired by the sparse coding hypothesis of visual cortex[olshausen1996emergence, barlow1961possible]:
@@
The dataset consists of 5,000 random $16\times 16$ grayscale patches extracted from CIFAR-10 images (converted from $32\times 32$ RGB via standard luminance formula)[krizhevsky2009learning]. Patches are normalized to the [0,1] range. We systematically compare eight input encoding strategies, which are grouped by their temporal characteristics[auge2021survey, guo2021neural]:
@@
Training uses the Adam optimizer[kingma2014adam] with a learning rate of $1\times 10^{-3}$ using a batch size of 64.
@@
  • Resolution: We repeated experiments on STL-10 ($96\times 96$ patches)[coates2011analysis]. Table 13 confirms the findings hold for different dataset and higher-resolution images.
Also, to confirm that the principle is not an artifact of the Duffing oscillator (Dynamical Universality), we replicated the receptive field learning experiment using two distinct systems: the Lorenz attractor (atmospheric convection)[lorenz2017deterministic] and the Thomas cyclically symmetric attractor (lattice dynamics)[thomas1999deterministic].
  • Lorenz System (Control for Entropy): Defined by $\dot{x}=\sigma(y-x)$, $\dot{y}=x(\rho-z)-y$, $\dot{z}=xy-\beta z$, with standard parameters $\sigma=10,\beta=8/3$. We scanned the Rayleigh number $\rho\in[0.1,28]$. In this setup, the phase space contraction rate remains constant ($\nabla\cdot F=-(\sigma+\beta+1)$) while $\rho$ increases the topological entropy.
  • Thomas Attractor (Control for Dissipation): Defined by $\dot{x}=\sin(y)-bx$, $\dot{y}=\sin(z)-by$, $\dot{z}=\sin(x)-bz$. The parameter $b$ linearly controls the global energy dissipation ($\nabla\cdot F=-3b$).
As illustrated in Figure 3, both systems exhibit a distinct phase transition dependent on their dynamical regimes. For the Lorenz system, the network successfully learns structured representations in the stable regime ($\rho<1.5$) but degrades into unstructured noise as the system enters the classical chaotic regime (e.g., $\rho=28$). This confirms that chaotic complexity alone without sufficient dissipative constraint is destructive. Similarly, the Thomas system demonstrates that at low damping ($b<0.2$, hyper-chaotic), the network fails to learn structure, whereas strong dissipation ($b\to 1.0$) forces the system into a constrained regime where structured features spontaneously emerge.
@@
Fine-Grained Landscape Validation. To rigorously decouple the effects of physical dynamics from explicit regularization, we conducted a high-resolution parameter sweep over the dynamical control parameter $\delta\in[-1.0,5.0]$ under varying sparsity weights ($\lambda$). Figure 4 reveals in-depth interaction patterns between dynamical complexity and learnability.
@@
Spectral Analysis and Scale-Space Invariance. To investigate the physical mechanism underlying the structural emergence, we performed a spectral analysis of the dynamical trajectories generated by the encoder. We estimated the Power Spectral Density (PSD) using Welch’s method[welch2003use] and extracted two key metrics: the Spectral Centroid, representing the weighted mean frequency of the signal, and the Spectral Entropy, quantifying the complexity of the power distribution.
Scale-Space Invariance (Fig. 5A). To verify that these spectral properties are intrinsic to the dynamical regime rather than artifacts of specific hyperparameters, we performed a multi-scale sweep across observation scales ($N$) and physical evolution times ($T_{max}\in\{4,8,12,16\}$). This analysis reveals a persistent spectral signature localized strictly within the transition regime, characterized by the stable coexistence of low dominant frequencies and high spectral entropy. Furthermore, this signature remains robust across varying timescales, confirming that transition dynamics actively suppress the accumulation of chaotic divergence. The sharp contrast between this ordered state and the high-frequency stochasticity of the expansive regime establishing it as a robust computational spot for generalization capacity.
@@ A.3.1 Encoding-Level Experiments
Setup. Agents were trained exclusively in an "Easy" environment (pole length 0.5m, gravity $9.8~m/s^{2}$) based on the CartPole-v1 task [towers2024gymnasium]. They were evaluated zero-shot on three physical properties to create progressively challenging environments while maintaining the same state-space structure, as presented in Table 14.
@@
These perturbations alter the system’s moment of inertia and responsiveness without changing the observation space or action set, providing a test of learned policies’ ability to generalize across physical parameter shifts[tobin2017domain].
All policy networks (SNN and MLP) shared an identical fully-connected topology of two hidden layers with 128 units each. This width was chosen to ensure sufficient network capacity, guaranteeing that performance differences would arise from the quality of the input encoding and architectural synergy, rather than from an under-parameterized model bottleneck. We use the REINFORCE algorithm[williams1992simple] with Adam optimizer ($lr=10^{-3}$, $\gamma=0.99$)[kingma2014adam]. All results are averaged over 10 independent runs.
@@
To ensure rigorous comparison, we incorporated Layer Normalization within the MLP hidden layers[ba2016layer]. This design choice was driven by our preliminary ablations: without extrinsic normalization, the stateless MLP suffered rapid numerical divergence when processing the high-variance Expansive ($\delta=-1.5$) encoding. Furthermore, while the Transition and Dissipative regimes achieved comparable mean performance without normalization, they exhibited significantly amplified inter-trial variance. We therefore standardized on Layer Normalization to eliminate optimization instability as a confounder, ensuring that performance differences reflect genuine representational capabilities rather than gradient issues. In contrast, the SNN required no such extrinsic stabilization; its intrinsic leaky-integration mechanism ($\beta=0.95$) provides natural temporal smoothing and dynamic stability.
@@ A.3.2 Architecture-Level Experiment Setup
The choice of RL algorithm was driven by task complexity. For CartPole (4D state, 200-step episodes), vanilla REINFORCE[williams1992simple] provided sufficient learning signal. For LunarLander (8D state, 1000-step episodes), the higher-dimensional state space and longer episodes required PPO[schulman2017proximal] with Generalized Advantage Estimation (GAE, $\lambda=0.95$) for stable training. Critically, neither algorithm introduces confounding temporal mechanisms: both remain policy gradient methods without learned world models, ensuring that performance differences reflect architectural properties rather than algorithmic artifacts.
@@
RLeaky SNNs provide a critical test case to disentangle these mechanisms. Unlike feedforward Leaky SNNs, which rely solely on membrane potential for temporal integration, RLeaky SNNs incorporate lateral recurrent connections that provide an additional memory pathway through persistent hidden states. This architecture increases the network’s capacity to retain information across timesteps[maass2002real].
@@ A.4 PAC-Bayesian Analysis of Dynamical Regimes
and are optimized via the reparameterization trick[blundell2015weight, kingma2013auto]. The network integrates spikes over $T=5$ simulation steps to produce logits, trained with cross-entropy loss. Training uses the Adam optimizer (lr $10^{-3}$), batch size $32$, and $100$ epochs per regime.
@@
summing over all network parameters. Given $\hat{L}_{\text{train}}$, training set size $m=m_{\text{train}}$, and confidence parameter $\delta_{\text{conf}}=0.05$, we evaluate the standard PAC-Bayesian upper bound[mcallester1998some] on the true risk $L(q)$, which holds with probability at least $1-\delta_{\text{conf}}$:
@@
The transition regime around $\delta=2.0$ attains the lowest $\mathrm{CV}_{\text{grad}}$ and the highest mean gradient magnitude $\mu_{grad}$, indicating a critical point that balances signal preservation and noise suppression. Expansive regimes exhibit large, highly variable gradients, consistent with the chaotic amplification of perturbations, whereas strongly dissipative regimes suppress gradients to the point that noise dominates, slightly increasing relative variability again.

### v2

@@ A.1 Experiment 1: Cross-Encoding Generalization Details
This appendix provides detailed experimental setups and supplementary analyses for the representational level experiments described in Section 2.2.
Dataset and Setup. We use the sklearn.datasets.load_digits dataset (64 features, 10 classes) [2, 45], split into 70% training, 15% validation, and 15% testing. Input features $\mathbf{x}$ are first normalized using StandardScaler. Networks are trained on data encoded with one $\delta_{\text{train}}$ value, then tested on 12 different encodings: $\delta_{\text{test}}\in\{-1.5,-1.0,-0.3,0.0,0.3,1.0,1.5,2.0,2.5,5.0,7.0,10.0\}$. The detailed quantitative generalization matrices (Mean $\pm$ Std %) for all architectures are provided in Tables 5 - 9.
Training. LIF neurons use learnable thresholds and membrane decay $\beta=0.95$. Training uses surrogate gradients (surrogate.fast_sigmoid with slope=25[19] with the Adam optimizer (lr=1e-4[29] for 200 epochs, with early stopping (patience=10) based on validation accuracy.
@@ A.2 Experiment 2: Unsupervised Receptive Field Learning Details
This appendix provides detailed experimental setups and supplementary analyses for the structural level experiments described in Section 2.3.
Dataset and Setup. The model for this experiment is a spiking neural network (SNN) autoencoder. Its architecture consists of an encoder linear layer ($W_{\text{enc}}\in\mathbb{R}^{128\times d}$), a recurrent Leaky Integrate-and-Fire (LIF) hidden layer with 128 learnable thresholds and membrane decay $\beta=0.95$ neurons [19], and a decoder linear layer ($W_{\text{dec}}\in\mathbb{R}^{d\times 128}$). The sparse autoencoder minimizes a composite loss balancing reconstruction fidelity and code sparsity, inspired by the sparse coding hypothesis of visual cortex [43, 7]:
@@
The dataset consists of 5,000 random $16\times 16$ grayscale patches extracted from CIFAR-10 images (converted from $32\times 32$ RGB via standard luminance formula) [33]. Patches are normalized to the [0,1] range. We systematically compare eight input encoding strategies, which are grouped by their temporal characteristics [4, 22]:
@@
Training uses the Adam optimizer [29] with a learning rate of $1\times 10^{-3}$ using a batch size of 64. The encoder weight matrix $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{128\times 256}$ maps flattened image patches to the bottleneck layer. Each row of $\mathbf{W}_{\text{enc}}$ represents the synaptic weights connecting all 256 input pixels to a single bottleneck neuron. To visualize the learned receptive field of neuron $i$, we reshape its weight vector $\mathbf{w}_{i}\in\mathbb{R}^{256}$ back to the original spatial dimensions ($16\times 16$). Figure 2,  3, and  4 display grids of such receptive fields, where red and blue pixels indicate positive (excitatory) and negative (inhibitory) connection weights, respectively. The standard deviation of these weights ($\sigma_{\text{RF}}$) quantifies structural differentiation: high variance indicates spatially organized filters, while low variance indicates homogeneous noise.
@@
  2. 2.
    Resolution: We repeated experiments on STL-10 ($96\times 96$ patches) [15]. Table 13 confirms the findings hold for a different dataset and higher-resolution images.
  3. 3.
    Dynamical Universality: To confirm that the principle is not an artifact of the duffing oscillator, we replicated the receptive field learning experiment using two distinct systems: the Lorenz attractor (atmospheric convection) [38] and the Thomas cyclically symmetric attractor (lattice dynamics) [53].
For the dynamical universality validation, we selected systems that allow independent control of dissipation versus dynamical complexity:
  • Lorenz System (varying entropy at fixed dissipation): Defined by $\dot{x}=\sigma(y-x)$, $\dot{y}=x(\rho-z)-y$, $\dot{z}=xy-\beta z$, with standard parameters $\sigma=10,\beta=8/3$. We scanned the Rayleigh number $\rho\in[0.1,28]$. In this setup, the phase space contraction rate remains constant ($\nabla\cdot F=-(\sigma+\beta+1)$) while $\rho$ increases the topological entropy.
  • Thomas Attractor (varying dissipation at fixed topology): Defined by $\dot{x}=\sin(y)-bx$, $\dot{y}=\sin(z)-by$, $\dot{z}=\sin(x)-bz$. The parameter $b$ linearly controls the global energy dissipation ($\nabla\cdot F=-3b$).
As illustrated in Figure 3, both systems exhibit a distinct phase transition dependent on their dynamical regimes: Lorenz system successfully learns structured representations in the stable regime ($\rho<1.5$) but degrades into unstructured noise as the system enters the classical chaotic regime (e.g., $\rho=28$). This confirms that chaotic complexity alone without sufficient dissipative constraint is destructive. Similarly, the Thomas system demonstrates that at low damping ($b<0.2$, hyper-chaotic), the network fails to learn structure, whereas strong dissipation ($b\to 1.0$) forces the system into a constrained regime where structured features spontaneously emerge.
@@
Fine-Grained Landscape Validation. To rigorously decouple the effects of dynamics from explicit regularization, we conducted a high-resolution parameter sweep over the dynamical control parameter $\delta\in[-1.0,5.0]$ under varying sparsity weights ($\lambda$). Figure 4 reveals in-depth interaction patterns between dynamical complexity and learnability.
@@
Spectral Analysis and Scale-Space Invariance. To investigate the physical mechanism underlying the structural emergence, we performed a spectral analysis of the dynamical trajectories generated by the encoder. We estimated the Power Spectral Density (PSD) using Welch’s method [57] and extracted two key metrics: the Spectral Centroid, representing the weighted mean frequency of the signal; and the Spectral Entropy, quantifying the complexity of the power distribution.
Scale-Space Invariance (Fig. 5A). To verify that these spectral properties are intrinsic to the dynamical regime rather than artifacts of specific hyperparameters, we performed a multi-scale sweep across observation scales ($N$) and physical evolution times ($T_{max}\in\{4,8,12,16\}$). This analysis reveals a persistent spectral signature localized strictly within the transition regime, characterized by the stable coexistence of low dominant frequencies and high spectral entropy. Furthermore, this signature remains robust across varying timescales, confirming that transition dynamics actively suppress the accumulation of chaotic divergence. The sharp contrast between this ordered state and the high-frequency stochasticity of the expansive regime establishes it as a robust computational spot for generalization capacity.
@@ A.3.1 Encoding-Level Experiments
Setup. Agents were trained exclusively in an "Easy" environment (pole length 0.5m, gravity $9.8~m/s^{2}$) based on the CartPole-v1 task [55]. They were evaluated zero-shot on three physical properties to create progressively challenging environments while maintaining the same state-space structure, as presented in Table 14.
@@
These perturbations alter the system’s moment of inertia and responsiveness without changing the observation space or action set, providing a test of learned policies’ ability to generalize across physical parameter shifts [54].
All policy networks (SNN and MLP) shared an identical fully-connected topology of two hidden layers with 128 units each. This width was chosen to ensure sufficient network capacity, guaranteeing that performance differences would arise from the quality of the input encoding and architectural synergy, rather than from an under-parameterized model bottleneck. We use the REINFORCE algorithm [58] with Adam optimizer ($lr=10^{-3}$, $\gamma=0.99$[29]. All results are averaged over 10 independent runs.
@@
To ensure rigorous comparison, we incorporated Layer Normalization within the MLP hidden layers [5]. This design choice was driven by our preliminary ablations: without extrinsic normalization, the stateless MLP suffered rapid numerical divergence when processing the high-variance Expansive ($\delta=-1.5$) encoding. Furthermore, while the Transition and Dissipative regimes achieved comparable mean performance without normalization, they exhibited significantly amplified inter-trial variance. We therefore standardized on Layer Normalization to eliminate optimization instability as a confounder, ensuring that performance differences reflect genuine representational capabilities rather than gradient issues. In contrast, the SNN required no such extrinsic stabilization; its intrinsic leaky-integration mechanism ($\beta=0.95$) provides natural temporal smoothing and dynamic stability.
@@ A.3.2 Architecture-Level Experiment Setup
The choice of RL algorithm was driven by task complexity. For CartPole (4D state, 200-step episodes), vanilla REINFORCE [58] provided sufficient learning signal. For LunarLander (8D state, 1000-step episodes), the higher-dimensional state space and longer episodes required PPO [51] with Generalized Advantage Estimation (GAE, $\lambda=0.95$) for stable training. Critically, neither algorithm introduces confounding temporal mechanisms: both remain policy gradient methods without learned world models, ensuring that performance differences reflect architectural properties rather than algorithmic artifacts.
@@
RLeaky SNNs provide a critical test case to disentangle these mechanisms. Unlike feedforward Leaky SNNs, which rely solely on membrane potential for temporal integration, RLeaky SNNs incorporate lateral recurrent connections that provide an additional memory pathway through persistent hidden states. This architecture increases the network’s capacity to retain information across timesteps [39].
@@ A.4 PAC-Bayesian Analysis of Dynamical Regimes
and are optimized via the reparameterization trick [10, 30]. The network integrates spikes over $5$ simulation steps to produce logits, trained with cross-entropy loss. Training uses the Adam optimizer (lr $10^{-3}$), batch size $32$, and $100$ epochs per regime.
@@
summing over all network parameters. Given $\hat{L}_{\text{train}}$, training set size $m=m_{\text{train}}$, and confidence parameter $\delta_{\text{conf}}=0.05$, we evaluate the standard PAC-Bayesian upper bound [40] on the true risk $L(q)$, which holds with probability at least $1-\delta_{\text{conf}}$:
@@
The transition regime around $\delta=2.0$ presents the lowest $\mathrm{CV}_{\text{grad}}$ and the highest mean gradient magnitude $\mu_{grad}$, indicating a critical point that balances signal preservation and noise suppression. Expansive regimes exhibit large, highly variable gradients, consistent with the chaotic amplification of perturbations, whereas strongly dissipative regimes suppress gradients to the point that noise dominates, slightly increasing relative variability again.